# HQ-SVC: Towards High-Quality Zero-Shot Singing Voice Conversion in Low-Resource Scenarios

### v1

@@ Introduction
Existing SVC methods (Liu et al., 2021b ; SVC-Develop-Team, 2023; Zhou et al., 2023; Bai et al., 2024) have demonstrated high-quality singing voice conversion by leveraging generative models like VITS (Kim et al., 2021) and diffusion (Ho et al., 2020) models. However, these models typically rely on explicit speaker IDs and separate content encoders (Hsu et al., 2021; Qian et al., 2022; Chen et al., 2022), which limit their generalization ability to unseen speakers, making them unsuitable for zero-shot singing voice conversion. FastSVC (Liu et al., 2021a ) achieves zero-shot singing voice conversion with a lightweight model but struggles with poor synthesis quality. LDM-SVC (Chen et al., 2024) provides high-quality zero-shot singing voice conversion on small datasets but demands significant computational resources and training time due to its two-stage and adversarial training. SaMoye-SVC (Wang et al., 2024b ) also delivers high-quality zero-shot singing voice conversion but requires large datasets and extensive adversarial training, making it highly resource-intensive. Moreover, these methods often rely on separate content and speaker encoders, leading to information loss and hindering effective feature fusion and reconstruction of natural and expressive voice.
@@
Given that both speech and singing belong to voice, the inherent similarities between them allow large-scale speech data to provide valuable priors, enabling transfer learning methods to effectively improve performance in zero-shot singing voice conversion tasks. Recent advancements in vector quantization (VQ) for zero-shot text-to-speech models like NaturalSpeech3 (Ju et al., 2024) and UniCATS (Du et al., 2024) have achieved significant success. NaturalSpeech3 introduces a VQ-based audio codec, FACodec, which turn audio into distinct latent spaces for content and speaker features simultaneously. This architecture allows zero-shot disentanglement of speaker and content features for unseen speakers, while also reducing information loss and making feature alignment easier compared to separate modeling. Although FACodec enables effective disentanglement of speaker and content features, it does not fully capture the complex acoustic variations necessary for high-quality synthesis, especially in challenging zero-shot singing voice conversion tasks.
@@
Therefore, based on unified decoupled audio codec, we propose HQ-SVC (High-Quality Zero-Shot Singing Voice Conversion), a novel framework designed to achieve high-quality zero-shot singing voice conversion in low-resource scenarios. Additionally, to improve high-quality synthesis modeling, we first leverage decoupled codec’s zero-shot disentanglement capability to extract content and speaker features from the decoder’s intermediate layer outputs. We then introduce supplementary features to improve the performance of SVC. To further enhance feature disentanglement and fusion, we propose the Enhanced Voice Adaptation (EVA) module, which incorporates additional acoustic features for multi-feature fusion, introduces a speaker loss, and includes a Speaker-F0 Predictor to enhance synthesis quality. Following (Yxlllc, 2023), the synthesis process is progressively optimized using Differentiable Digital Signal Processing (DDSP) (Engel et al., 2020) and diffusion (Ho et al., 2020) models to generate Mel spectrograms. Finally, a vocoder is employed to produce singing voices.
@@
Singing Voice Conversion (SVC) aims to transform a singer’s timbre into that of another target singer while preserving the melody and content. Early SVC methods were primarily based on Gaussian Mixture Models (GMM) (Doi et al., 2012; Kobayashi et al., 2014, 2015), which relied on parallel data. These methods require both source and target singers to perform the same song. This dependency not only demanded a large amount of parallel data but also limited the generation quality of the models. With the rise of Generative Adversarial Networks (GANs) (Goodfellow et al., 2020), several GAN-based SVC approaches (Sisman et al., 2019; Sisman and Li, 2020; Liu et al., 2021a ) emerged that operate without parallel data. These methods employ adversarial training, where a generator and a discriminator engage in a game to improve the audio quality. However, GAN-based methods typically take longer to train and can suffer from issues like mode collapse, making the overall training process more challenging. SoVITS-SVC (SVC-Develop-Team, 2023) builds on adversarial training by adding a flow-based model to prevent mode collapse during training. It is based on VITS (Kim et al., 2021), which uses an end-to-end flow-based architecture with adversarial training to generate high-quality singing voices. Despite its effectiveness, it still demands substantial computational resources for training.
@@
Open-source singing voice datasets available for academic research remain relatively scarce. At the same time, the growing practical demand for efficient training and inference on consumer-grade GPUs has led to increasing research attention on achieving high-quality SVC under low-resource conditions. Low-resource SVC means using only a small amount of singing voice data and limited GPU power during training, while still producing high-quality audio. Several methods have been proposed to address this challenge. These include DDSP-SVC (Yxlllc, 2023) based on Differentiable Digital Signal Processing (DDSP) models (Engel et al., 2020) and diffusion-based SVC methods(Liu et al., 2021b ; Lu et al., 2024; Huang et al., 2025). These methods not only demonstrate superior generation quality, but also exhibit strong training stability, enabling high-quality singing voice conversion even in low-resource scenarios.
@@
Nevertheless, while the aforementioned SVC methods have made significant strides, their reliance on explicit speaker identities restricts their ability to generalize to unseen speakers, making them unsuitable for zero-shot SVC scenarios. In zero-shot SVC, the objective is to achieve high-quality conversion without requiring target singer data or fine-tuning. Traditional SVC approaches typically depend on parallel data and explicit speaker labels, and in practical applications, issues such as high sampling rates, broader pitch ranges, and the intricate coupling between timbre and pitch present significant challenges in modeling and training. FastSVC (Liu et al., 2021a ) is one of the earlier zero-shot singing voice conversion methods; it leverages GANs combined with FiLM (Perez et al., 2018) to enable a lightweight zero-shot conversion. However, its generation quality leaves room for improvement. In contrast, methods like LDM-SVC (Chen et al., 2024), FreeSVC (Ferreira et al., 2025), and SaMoye-SVC (Wang et al., 2024b ), which are based on VITS, can synthesize high-quality singing voices in a zero-shot manner, although at the cost of requiring considerable computational power during training.
@@ Methods
We utilize FACodec (Ju et al., 2024) as the unified disentangler. Specifically, we freeze all of the layers of the encoder and decoder of the FACodec and get the $x_{\mathrm{con}}$ and $x_{\mathrm{spk}}$ as the content and speaker features respectively. The basic architecture of FACodec follows Kumar et al. (Kumar et al., 2024) and employs the SnakeBeta activation function (Lee et al., 2023). The speaker extractor consists of several Conformer (Gulati et al., 2020) blocks. For the three factorized vector quantization (FVQ) components $Q^{c}$, $Q^{p}$, $Q^{d}$, we use $N_{q_{c}}$ = 2, $N_{q_{p}}$ = 1, $N_{q_{d}}$ = 3 quantizers respectively, with all quantizers having a codebook size of 1024.
@@
DDSP (Engel et al., 2020) is a method that integrates classical digital signal processing (DSP) techniques with deep learning. It introduces strong inductive biases while preserving the expressive power of neural networks, enabling high-fidelity audio synthesis without relying on large autoregressive models or adversarial losses. Its modular architecture allows for fine-grained control over attributes such as timbre and loudness. Specifically, the DDSP synthesizer consists of a harmonic synthesizer and a noise synthesizer, which generate periodic and aperiodic components respectively and combine them to produce the final waveform. Incorporating DDSP into the SVC task enhances control over audio attributes, thereby improving conversion performance.
@@
Moreover, to improve the quality of vocal synthesis, the diffusion (Ho et al., 2020) model is utilized to process the output of DDSP. Integrating the diffusion model with DDSP enriches the description of sound characteristics by filling in any gaps, thereby boosting the performance of SVC. The strong performance of diffusion model in SVC tasks has been demonstrated in DiffSVC (Liu et al., 2021b ) and CoMoSVC (Lu et al., 2024).
@@
Singing voices are often rich in melody and energy dynamics, making it challenging to ensure high-quality generation using only content and speaker features. To address this, we propose the Enhanced Voice Adaptation (EVA) module, which introduces RMVPE (Wei et al., 2023) as an F0 encoder to extract pitch features and incorporates energy features along with phase features required by the DDSP (Engel et al., 2020) model for multi-feature fusion.
@@
As shown in Fig. 1 (b), since FACodec’s speaker features $x_{\mathrm{spk}}$ may include unrelated characteristics like style, they are mapped to speaker embeddings $e_{\mathrm{spk}}$ using an MLP. And the residual between the speaker features $x_{\mathrm{spk}}$ and speaker embeddings $e_{\mathrm{spk}}$ is computed to obtain residual style features $x_{\mathrm{sty}}$, which is expanded to get $e_{\mathrm{sty}}$. Following the approach in DDSP-SVC (Yxlllc, 2023), we perform the transformation on the F0 to obtain
@@
to help the model better capture the melody. And we map it to pitch embeddings $e_{f_{0}}$ through an MLP. Similarly, volume feature $x_{\mathrm{vol}}$ and phase $x_{\mathrm{pha}}$ feature are fed into MLPs to generate their respective embeddings $e_{\mathrm{vol}}$ and $e_{\mathrm{pha}}$. All the above MLPs consist of two linear layers, with a SiLU (Elfwing et al., 2018) activation function applied between them. The output dimension of MLPs is 256.
@@
which is subsequently compressed to 256 dimensions via a 1D convolution to match the content. The content features $x_{\mathrm{con}}$, after convolution and normalization, are fused with the aforementioned style embeddings using FiLM (Perez et al., 2018), as described by
@@
where $e_{c}$ denotes the content embeddings, $e_{s}$ represents the style embeddings, and $f_{\alpha}(\cdot)$ and $f_{\beta}(\cdot)$ correspond to two independent linear layers. To achieve more effective fusion of features, the FiLM output is passed through a Conformer (Gulati et al., 2020; Vaswani, 2017) with 8-head self-attention and layer normalization, producing $e$, which is then used as input for the DDSP and diffusion models.
In addition, to enhance the model’s performance in zero-shot singing voice conversion for unseen speakers, we introduce the speaker loss $\mathcal{L}_{\mathrm{spk}}$ based on the InfoNCE loss (Oord et al., 2018), a widely-used contrastive learning objective. It helps the model distinguish speakers by pulling positive samples closer and pushing negative samples farther within a batch:
@@
Additionally, we pass $e$ to the diffusion model for further optimization. We use the WaveNet (Oord et al., 2016) from FastSpeech2 (Ren et al., 2020) as the diffusion denoiser, configured with 128-dimensional input features, 20 residual block layers, 512 convolutional output channels, and a 256-dimensional encoder hidden layer. The formula for the diffusion loss is given by:
@@
During inference, DPM-Solver++ (Lu et al., 2022) is used for sampling in the diffusion model, with chunk-wise processing adopted to reduce inference latency, 100 diffusion steps and a speed-up rate of 10 to balance quality and speed. To improve synthesis quality, HQ-SVC uses NSF-HiFiGAN (OpenVPI, 2022), an enhanced version of HiFi-GAN (Kong et al., 2020) with a Neural Source-Filter (NSF) (Wang et al., 2019) module, to convert Mel spectrogram with F0 into audio.
@@
In this study, we utilized the high-quality, open-source Mandarin singing datasets Opensinger (Huang et al., 2021) and M4Singer (Zhang et al., 2022b ) for model training. From Opensinger, 2 male and 2 female singers were randomly selected as unseen singers. These singers were excluded from both training and validation and were only used for model testing. Additionally, 1% of the remaining singing data from M4Singer and Opensinger was randomly selected as a validation set to monitor model convergence. To avoid potential noise and performance degradation caused by excessively short audio samples, we excluded audio clips shorter than 2.1 seconds.
@@
To enable zero-shot singing voice conversion in low-resource scenarios, we utilized frozen pre-trained models for feature extraction, effectively reducing memory consumption during training and inference. Since the singing audio files have two different rates: 44.1 kHz and 48 kHz, we resampled them to 44.1 kHz. Then, we used a hop size of 512 and obtained 128-dimensional Mel spectrograms and energy features from the audios. Next, we downsampled the audios to 16 kHz, used RMVPE to extract pitch features, and employed the FACodec (Ju et al., 2024; Zhang et al., 2024) to extract 256-dimensional content and speaker features. Furthermore, we constructed a lookup table to map all speakers across datasets to unique speaker IDs in a unified table for speaker loss calculation.
@@
During training, we employed the AdamW (Loshchilov and Hutter, 2017) optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, an initial learning rate of $1.5\times 10^{-4}$, a batch size of $N=64$. For the speaker loss, the temperature parameter was set to $\tau=0.1$. The model was trained for 250k steps on a single NVIDIA RTX 3090 GPU. Notably, HQ-SVC occupies less than 6 GB of GPU memory during training and can be trained within about 11 hours, making it well-suited for zero-shot singing voice conversion in low-resource scenarios.
@@
We conducted a comparative evaluation of the proposed HQ-SVC, against the SOTA zero-shot singing voice conversion model. Since most zero-shot SVC methods are not open-source, the VITS-based SaMoye-SVC (Wang et al., 2024b ) model pretrained on 1700 hours of clean singing dataset, was used as a strong baseline for comparison, whereas HQ-SVC was trained on less than 80 hours of singing data. To further validate the effectiveness of our improvements to the FACodec, we fine-tuned FACodec. For fairness, we modified FACodec’s output structure to produce Mel spectrograms, which were combined with F0 and synthesized into singing audio using NSF-HiFiGAN (OpenVPI, 2022), like HQ-SVC. In addition, all generated audios were resampled to 32 kHz for zero-shot singing voice conversion evaluation, since HQ-SVC generates audios at 44.1 kHz, while SaMoye outputs at 32 kHz.
We assessed the models on unseen speakers using objective measures—intelligibility (STOI), timbre similarity (SECS), pitch accuracy (F0 RMSE and FPC), overall quality (NISQA) (Mittag et al., 2021) and subjective 5-point ratings of naturalness (NMOS) and speaker similarity (SMOS).
@@
HQ-SVC utilizes 16 kHz data for feature extraction during training and employs authentic 44.1 kHz Mel spectrograms for model optimization. This design allows HQ-SVC to achieve zero-shot voice super-resolution (SR) using the zero-shot singing voice conversion method while preserving the speaker’s identity. To demonstrate this, we compared HQ-SVC with the current SOTA method AudioSR (Liu et al., 2024). AudioSR uses a Latent Diffusion Model (LDM) trained on over 7,000 hours of sound effects, speech, and music, allowing it to upscale audio to 48 kHz.
Since HQ-SVC uniformly resamples ground-truth audio to 44.1 kHz during training, we set its output sampling rate to 44.1 kHz for a fair comparison. We ensure fairness and demonstrate the generalization of HQ-SVC on Voice Super-resolution tasks for both singing and speech using unseen test datasets. These include Mandarin Opensinger and English NHSS-Song (Sharma et al., 2021) (singing datasets), as well as English LibriTTS-P (Kawamura et al., 2024) and NHSS-Speech (speech datasets).
For evaluation on voice super-resolution task, we randomly selected 2 male and 2 female speakers from each of the four datasets, totaling 8 male and 8 female speakers. From each speaker, we randomly chose 8 audio samples, giving us 128 samples for objective evaluation. For subjective evaluation, we randomly selected 1 audio sample per speaker, resulting in 8 singing and 8 speech samples. This yielded a total of 96 singing and 96 speech samples per model. The objective metrics include STOI, SECS, FPC, NISQA, and log-spectral distance (LSD) (Gray and Markel, 1976) metric adopted by AudioSR, while the subjective metrics include NMOS and SMOS.
@@
However, we noticed that while SaMoye performs slightly worse in SMOS, it performs better in the SECS metric. This difference may be due to how SECS and human perception respond to pitch. SECS is calculated using cosine similarity from speaker embeddings extracted by the CAM++ (Wang et al., 2023a ) speaker verification model. These embeddings implicitly contain statistical information about pitch, allowing model to effectively identify the speaker in the converted audio, even when the pitch remains unchanged. In contrast, human perception of speaker identity is more sensitive to variations in pitch. Thus, even if the speaker characteristics are altered, a consistent pitch can make it difficult for humans to identify the speaker.
The results of the ablation experiments further validate the above observations. Removing either the $\mathcal{L}_{f_{0}}$ or the $\mathcal{L}_{\mathrm{spk}}$ leads to a varying degrees of decline in SECS, F0 RMSE, NISQA, and NMOS. This indicates that these two losses contribute to generating singing voices with more accurate pitch, while ensuring that the synthesized voices remain natural and are more easily distinguishable by the speaker verification model. We also observed a slight improvement in SMOS when $\mathcal{L}_{f_{0}}$ or $\mathcal{L}_{\mathrm{spk}}$ is removed. This further supports the earlier point: adding $\mathcal{L}_{f_{0}}$ and $\mathcal{L}_{\mathrm{spk}}$ enhances pitch-related information in the speaker embeddings, making them more identifiable by CAM++ (Wang et al., 2023a ) and better at maintaining a consistent F0. This consistency makes it difficult for human listeners to perceive changes in speaker characteristics, particularly when the target speaker has a significantly different vocal range, thus leading to a decrease in SMOS. Nevertheless, the timbre similarity of HQ-SVC remains superior to other baseline models.
@@
Table 4 presents a comparison of zero-shot singing voice conversion performance under various acceleration rates (1×, 2×, 10×, 20×) using three diffusion-based samplers: Dpm-solver++ (Lu et al., 2022), UniPC (Zhao et al., 2023), and DDIM (Song et al., 2020).

### v2

@@ Introduction
Existing SVC methods (liu2021diffsvc; so-vits-svc2023; zhou2023vitssvcc; bai24_spa_svc) have demonstrated high-quality singing voice conversion by leveraging generative models like VITS (kim2021conditional_vits) and diffusion (ho2020ddpm) models. However, these models typically rely on explicit speaker IDs and separate content encoders (hsu2021hubert; qian2022contentvec; chen2022wavlm), which limit their generalization ability to unseen speakers, making them unsuitable for zero-shot singing voice conversion. FastSVC (liu2021fastsvc_cross) achieves zero-shot singing voice conversion with a lightweight model but struggles with poor synthesis quality. LDM-SVC (chen2024ldm) provides high-quality zero-shot singing voice conversion on small datasets but demands significant computational resources and training time due to its two-stage and adversarial training. SaMoye-SVC (wang2024samoye) also delivers high-quality zero-shot singing voice conversion but requires large datasets and extensive adversarial training, making it highly resource-intensive. Moreover, these methods often rely on separate content and speaker encoders, leading to information loss and hindering effective feature fusion and reconstruction of natural and expressive voice.
@@
Given that both speech and singing belong to voice, the inherent similarities between them allow large-scale speech data to provide valuable priors, enabling transfer learning methods to effectively improve performance in zero-shot singing voice conversion tasks. Recent advancements in vector quantization (VQ) for zero-shot text-to-speech models like NaturalSpeech3 (junaturalspeech3) and UniCATS (du2024unicats) have achieved significant success. NaturalSpeech3 introduces a VQ-based audio codec, FACodec, which turn audio into distinct latent spaces for content and speaker features simultaneously. This architecture allows zero-shot disentanglement of speaker and content features for unseen speakers, while also reducing information loss and making feature alignment easier compared to separate modeling. Although FACodec enables effective disentanglement of speaker and content features, it does not fully capture the complex acoustic variations necessary for high-quality synthesis, especially in challenging zero-shot singing voice conversion tasks.
@@
Therefore, based on unified decoupled audio codec, we propose HQ-SVC (High-Quality Zero-Shot Singing Voice Conversion), a novel framework designed to achieve high-quality zero-shot singing voice conversion in low-resource scenarios. Additionally, to improve high-quality synthesis modeling, we first leverage decoupled codec’s zero-shot disentanglement capability to extract content and speaker features from the decoder’s intermediate layer outputs. We then introduce supplementary features to improve the performance of SVC. To further enhance feature disentanglement and fusion, we propose the Enhanced Voice Adaptation (EVA) module, which incorporates additional acoustic features for multi-feature fusion, introduces a speaker loss, and includes a Speaker-F0 Predictor to enhance synthesis quality. Following (ddsp-svc2023), the synthesis process is progressively optimized using Differentiable Digital Signal Processing (DDSP) (engel2020ddsp) and diffusion (ho2020ddpm) models to generate Mel spectrograms. Finally, a vocoder is employed to produce singing voices.
@@
Singing Voice Conversion (SVC) aims to transform a singer’s timbre into that of another target singer while preserving the melody and content. Early SVC methods were primarily based on Gaussian Mixture Models (GMM) (doi2012gmmsvc; kobayashi2014gmmsvc; kobayashi2015gmmsvc), which relied on parallel data. These methods require both source and target singers to perform the same song. This dependency not only demanded a large amount of parallel data but also limited the generation quality of the models. With the rise of Generative Adversarial Networks (GANs) (goodfellow2020gan), several GAN-based SVC approaches (sisman2019singan; sisman2020generative; liu2021fastsvc_cross) emerged that operate without parallel data. These methods employ adversarial training, where a generator and a discriminator engage in a game to improve the audio quality. However, GAN-based methods typically take longer to train and can suffer from issues like mode collapse, making the overall training process more challenging. SoVITS-SVC (so-vits-svc2023) builds on adversarial training by adding a flow-based model to prevent mode collapse during training. It is based on VITS (kim2021conditional_vits), which uses an end-to-end flow-based architecture with adversarial training to generate high-quality singing voices. Despite its effectiveness, it still demands substantial computational resources for training.
@@
Open-source singing voice datasets available for academic research remain relatively scarce. At the same time, the growing practical demand for efficient training and inference on consumer-grade GPUs has led to increasing research attention on achieving high-quality SVC under low-resource conditions. Low-resource SVC means using only a small amount of singing voice data and limited GPU power during training, while still producing high-quality audio. Several methods have been proposed to address this challenge. These include DDSP-SVC (ddsp-svc2023) based on Differentiable Digital Signal Processing (DDSP) models (engel2020ddsp) and diffusion-based SVC methods(liu2021diffsvc; lu2024comosvc; huang2025lhq). These methods not only demonstrate superior generation quality, but also exhibit strong training stability, enabling high-quality singing voice conversion even in low-resource scenarios.
@@
Nevertheless, while the aforementioned SVC methods have made significant strides, their reliance on explicit speaker identities restricts their ability to generalize to unseen speakers, making them unsuitable for zero-shot SVC scenarios. In zero-shot SVC, the objective is to achieve high-quality conversion without requiring target singer data or fine-tuning. Traditional SVC approaches typically depend on parallel data and explicit speaker labels, and in practical applications, issues such as high sampling rates, broader pitch ranges, and the intricate coupling between timbre and pitch present significant challenges in modeling and training. FastSVC (liu2021fastsvc_cross) is one of the earlier zero-shot singing voice conversion methods; it leverages GANs combined with FiLM (perez2018film) to enable a lightweight zero-shot conversion. However, its generation quality leaves room for improvement. In contrast, methods like LDM-SVC (chen2024ldm), FreeSVC (ferreira2025freesvc), and SaMoye-SVC (wang2024samoye), which are based on VITS, can synthesize high-quality singing voices in a zero-shot manner, although at the cost of requiring considerable computational power during training.
@@ Methods
We utilize FACodec (junaturalspeech3) as the unified disentangler. Specifically, we freeze all of the layers of the encoder and decoder of the FACodec and get the $x_{\mathrm{con}}$ and $x_{\mathrm{spk}}$ as the content and speaker features respectively. The basic architecture of FACodec follows Kumar et al. (kumar2024_improved_rvqgan) and employs the SnakeBeta activation function (lee2023bigvgan). The speaker extractor consists of several Conformer (gulati20_interspeech_conformer) blocks. For the three factorized vector quantization (FVQ) components $Q^{c}$, $Q^{p}$, $Q^{d}$, we use $N_{q_{c}}$ = 2, $N_{q_{p}}$ = 1, $N_{q_{d}}$ = 3 quantizers respectively, with all quantizers having a codebook size of 1024.
@@
DDSP (engel2020ddsp) is a method that integrates classical digital signal processing (DSP) techniques with deep learning. It introduces strong inductive biases while preserving the expressive power of neural networks, enabling high-fidelity audio synthesis without relying on large autoregressive models or adversarial losses. Its modular architecture allows for fine-grained control over attributes such as timbre and loudness. Specifically, the DDSP synthesizer consists of a harmonic synthesizer and a noise synthesizer, which generate periodic and aperiodic components respectively and combine them to produce the final waveform. Incorporating DDSP into the SVC task enhances control over audio attributes, thereby improving conversion performance.
@@
Moreover, to improve the quality of vocal synthesis, the diffusion (ho2020ddpm) model is utilized to process the output of DDSP. Integrating the diffusion model with DDSP enriches the description of sound characteristics by filling in any gaps, thereby boosting the performance of SVC. The strong performance of diffusion model in SVC tasks has been demonstrated in DiffSVC (liu2021diffsvc) and CoMoSVC (lu2024comosvc).
@@
Singing voices are often rich in melody and energy dynamics, making it challenging to ensure high-quality generation using only content and speaker features. To address this, we propose the Enhanced Voice Adaptation (EVA) module, which introduces RMVPE (wei2023rmvpe) as an F0 encoder to extract pitch features and incorporates energy features along with phase features required by the DDSP (engel2020ddsp) model for multi-feature fusion.
@@
As shown in Fig. 1 (b), since FACodec’s speaker features $x_{\mathrm{spk}}$ may include unrelated characteristics like style, they are mapped to speaker embeddings $e_{\mathrm{spk}}$ using an MLP. And the residual between the speaker features $x_{\mathrm{spk}}$ and speaker embeddings $e_{\mathrm{spk}}$ is computed to obtain residual style features $x_{\mathrm{sty}}$, which is expanded to get $e_{\mathrm{sty}}$. Following the approach in DDSP-SVC (ddsp-svc2023), we perform the transformation on the F0 to obtain
@@
to help the model better capture the melody. And we map it to pitch embeddings $e_{f_{0}}$ through an MLP. Similarly, volume feature $x_{\mathrm{vol}}$ and phase $x_{\mathrm{pha}}$ feature are fed into MLPs to generate their respective embeddings $e_{\mathrm{vol}}$ and $e_{\mathrm{pha}}$. All the above MLPs consist of two linear layers, with a SiLU (elfwing2018silu) activation function applied between them. The output dimension of MLPs is 256.
@@
which is subsequently compressed to 256 dimensions via a 1D convolution to match the content. The content features $x_{\mathrm{con}}$, after convolution and normalization, are fused with the aforementioned style embeddings using FiLM (perez2018film), as described by
@@
where $e_{c}$ denotes the content embeddings, $e_{s}$ represents the style embeddings, and $f_{\alpha}(\cdot)$ and $f_{\beta}(\cdot)$ correspond to two independent linear layers. To achieve more effective fusion of features, the FiLM output is passed through a Conformer (gulati20_interspeech_conformer; vaswani2017attention) with 8-head self-attention and layer normalization, producing $e$, which is then used as input for the DDSP and diffusion models.
In addition, to enhance the model’s performance in zero-shot singing voice conversion for unseen speakers, we introduce the speaker loss $\mathcal{L}_{\mathrm{spk}}$ based on the InfoNCE loss (oord2018infonce), a widely-used contrastive learning objective. It helps the model distinguish speakers by pulling positive samples closer and pushing negative samples farther within a batch:
@@
Additionally, we pass $e$ to the diffusion model for further optimization. We use the WaveNet (oord2016wavenet) from FastSpeech2 (ren2020fastspeech) as the diffusion denoiser, configured with 128-dimensional input features, 20 residual block layers, 512 convolutional output channels, and a 256-dimensional encoder hidden layer. The formula for the diffusion loss is given by:
@@
During inference, DPM-Solver++ (lu2022dpm) is used for sampling in the diffusion model, with chunk-wise processing adopted to reduce inference latency, 100 diffusion steps and a speed-up rate of 10 to balance quality and speed. To improve synthesis quality, HQ-SVC uses NSF-HiFiGAN (openvpi2022nsf-hifigan), an enhanced version of HiFi-GAN (kong2020hifigan) with a Neural Source-Filter (NSF) (wang2019nsf) module, to convert Mel spectrogram with F0 into audio.
@@
In this study, we utilized the high-quality, open-source Mandarin singing datasets Opensinger (huang2021opensinger) and M4Singer (zhang2022m4singer) for model training. From Opensinger, 2 male and 2 female singers were randomly selected as unseen singers. These singers were excluded from both training and validation and were only used for model testing. Additionally, 1% of the remaining singing data from M4Singer and Opensinger was randomly selected as a validation set to monitor model convergence. To avoid potential noise and performance degradation caused by excessively short audio samples, we excluded audio clips shorter than 2.1 seconds.
@@
To enable zero-shot singing voice conversion in low-resource scenarios, we utilized frozen pre-trained models for feature extraction, effectively reducing memory consumption during training and inference. Since the singing audio files have two different rates: 44.1 kHz and 48 kHz, we resampled them to 44.1 kHz. Then, we used a hop size of 512 and obtained 128-dimensional Mel spectrograms and energy features from the audios. Next, we downsampled the audios to 16 kHz, used RMVPE to extract pitch features, and employed the FACodec (junaturalspeech3; amphion) to extract 256-dimensional content and speaker features. Furthermore, we constructed a lookup table to map all speakers across datasets to unique speaker IDs in a unified table for speaker loss calculation.
@@
During training, we employed the AdamW (loshchilov2017adamw) optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, an initial learning rate of $1.5\times 10^{-4}$, a batch size of $N=64$. For the speaker loss, the temperature parameter was set to $\tau=0.1$. The model was trained for 250k steps on a single NVIDIA RTX 3090 GPU. Notably, HQ-SVC occupies less than 6 GB of GPU memory during training and can be trained within about 11 hours, making it well-suited for zero-shot singing voice conversion in low-resource scenarios.
@@
We conducted a comparative evaluation of the proposed HQ-SVC, against the SOTA zero-shot singing voice conversion model. Since most zero-shot SVC methods are not open-source, the VITS-based SaMoye-SVC (wang2024samoye) model pretrained on 1700 hours of clean singing dataset, was used as a strong baseline for comparison, whereas HQ-SVC was trained on less than 80 hours of singing data. To further validate the effectiveness of our improvements to the FACodec, we fine-tuned FACodec. For fairness, we modified FACodec’s output structure to produce Mel spectrograms, which were combined with F0 and synthesized into singing audio using NSF-HiFiGAN (openvpi2022nsf-hifigan), like HQ-SVC. In addition, all generated audios were resampled to 32 kHz for zero-shot singing voice conversion evaluation, since HQ-SVC generates audios at 44.1 kHz, while SaMoye outputs at 32 kHz.
We assessed the models on unseen speakers using objective measures—intelligibility (STOI), timbre similarity (SECS), pitch accuracy (F0 RMSE and FPC), overall quality (NISQA) (mittag2021nisqa) and subjective 5-point ratings of naturalness (NMOS) and speaker similarity (SMOS).
@@
HQ-SVC utilizes 16 kHz data for feature extraction during training and employs authentic 44.1 kHz Mel spectrograms for model optimization. This design allows HQ-SVC to achieve zero-shot voice super-resolution (SR) using the zero-shot singing voice conversion method while preserving the speaker’s identity. To demonstrate this, we compared HQ-SVC with the current SOTA method AudioSR (liu2024audiosr). AudioSR uses a Latent Diffusion Model (LDM) trained on over 7,000 hours of sound effects, speech, and music, allowing it to upscale audio to 48 kHz.
Since HQ-SVC uniformly resamples ground-truth audio to 44.1 kHz during training, we set its output sampling rate to 44.1 kHz for a fair comparison. We ensure fairness and demonstrate the generalization of HQ-SVC on Voice Super-resolution tasks for both singing and speech using unseen test datasets. These include Mandarin Opensinger and English NHSS-Song (sharma2021nhss) (singing datasets), as well as English LibriTTS-P (kawamura24_libritts_p) and NHSS-Speech (speech datasets).
For evaluation on voice super-resolution task, we randomly selected 2 male and 2 female speakers from each of the four datasets, totaling 8 male and 8 female speakers. From each speaker, we randomly chose 8 audio samples, giving us 128 samples for objective evaluation. For subjective evaluation, we randomly selected 1 audio sample per speaker, resulting in 8 singing and 8 speech samples. This yielded a total of 96 singing and 96 speech samples per model. The objective metrics include STOI, SECS, FPC, NISQA, and log-spectral distance (LSD) (1976_lsd_distance_measures) metric adopted by AudioSR, while the subjective metrics include NMOS and SMOS.
@@
However, we noticed that while SaMoye performs slightly worse in SMOS, it performs better in the SECS metric. This difference may be due to how SECS and human perception respond to pitch. SECS is calculated using cosine similarity from speaker embeddings extracted by the CAM++ (wang23ha_interspeech_campp) speaker verification model. These embeddings implicitly contain statistical information about pitch, allowing model to effectively identify the speaker in the converted audio, even when the pitch remains unchanged. In contrast, human perception of speaker identity is more sensitive to variations in pitch. Thus, even if the speaker characteristics are altered, a consistent pitch can make it difficult for humans to identify the speaker.
The results of the ablation experiments further validate the above observations. Removing either the $\mathcal{L}_{f_{0}}$ or the $\mathcal{L}_{\mathrm{spk}}$ leads to a varying degrees of decline in SECS, F0 RMSE, NISQA, and NMOS. This indicates that these two losses contribute to generating singing voices with more accurate pitch, while ensuring that the synthesized voices remain natural and are more easily distinguishable by the speaker verification model. We also observed a slight improvement in SMOS when $\mathcal{L}_{f_{0}}$ or $\mathcal{L}_{\mathrm{spk}}$ is removed. This further supports the earlier point: adding $\mathcal{L}_{f_{0}}$ and $\mathcal{L}_{\mathrm{spk}}$ enhances pitch-related information in the speaker embeddings, making them more identifiable by CAM++ (wang23ha_interspeech_campp) and better at maintaining a consistent F0. This consistency makes it difficult for human listeners to perceive changes in speaker characteristics, particularly when the target speaker has a significantly different vocal range, thus leading to a decrease in SMOS. Nevertheless, the timbre similarity of HQ-SVC remains superior to other baseline models.
@@
Table 4 presents a comparison of zero-shot singing voice conversion performance under various acceleration rates (1×, 2×, 10×, 20×) using three diffusion-based samplers: Dpm-solver++ (lu2022dpm), UniPC (zhao2023unipc), and DDIM (song2020ddim).